# Bridging Operator Learning and Conditioned Neural Fields: A Unifying Perspective

### v1

@@ C.1 Training and evaluation
We use a unified training recipe for all CViT experiments. We employ AdamW optimizer [70, 71] with a weight decay $10^{-5}$. Our learning rate schedule includes an initial linear warm-up phase of $5,000$ steps, starting from zero and gradually increasing to $10^{-3}$, followed by an exponential decay at a rate of $0.9$ for every $5,000$ steps. The loss function is a one-step mean squared error (MSE) between the model predictions and the corresponding targets at the next time-step, evaluated at randomly sampled query coordinates:
@@
where $s^{(k)}_{i}(\mathbf{y_{j}})$ denotes the $k$-th variable of the $i$-th sample in the training dataset, evaluated at a query coordinate $\mathbf{y}_{j}$, and $\hat{s}$ denotes the corresponding model prediction. All models are trained for $2\times 10^{5}$ iterations with a batch size $B=64$. Within each batch, we randomly sample $Q=1,024$ query coordinates from the grid and corresponding output labels.
@@
After training, we obtain the predicted trajectory by performing an auto-regressive rollout on the test dataset. We evaluate model accuracy using the relative $L^{2}$ norm, as commonly used in Li et al.[4]:
@@ C.2 Model Details
Table 2: Details of Continuous Vision Transformer model variants.
Model Encoder layers Embedding dim MLP width Heads # Params
CViT-S 5 384 384 6 13 M
CViT-B 10 512 512 8 30 M
CViT-L 15 768 1536 12 92 M
@@
Following the comprehensive experiments and ablation studies in PDEArena [57], we adopt the optimal Dilated ResNet configuration reported in their work. This model has 128 channels and employs group normalization. It consists of four residual blocks, each comprising seven dilated CNN layers with dilation rates of [1, 2, 4, 8, 4, 2, 1].
@@
Following the comprehensive experiments and ablation studies in PDEArena [57], we adopt the optimal UNO of 128 channels.
@@
For the shallow water equation, we employ the best FNO model from comprehensive experiments and ablation studies in PDEArena [57]. Specifically, we use the FNO consisting of 4 FNO layers, each with 128 channels and 32 modes. Additionally, we use two embedding layers and two output layers with kernel sizes of $1\times 1$, as suggested in Li et al.[4]. We use GeLU activation functions and no normalization scheme.
For the Navier-Stokes equation, we follow the problem setup in DPOT [60], which differs from the setup in PDEArena [57]. Therefore, we directly report the results from DPOT [60].
FFNO / GK-T / OFormer / GNOT / DPOT.
We directly report the results from DPOT [60]. Detailed implementations can be found in their paper.
C.3 Linear Advection Equation
@@
We make use of the datasets released by Hoop et al.[61]. This dataset consists of 40,000 samples of discretized initial conditions on a grid of $N=200$ points.
@@
We follow the problem setup by Hoop et al.[61]. Our objective is to predict the solution profile at time $t=0.5$. We compare the CViT model’s performance against several strong baseline neural operators: DeepONet [6], NoMaD [12] and Fourier Neural Operators [4]. For the training and evaluation, we considered a split of 20,000 samples used for training, 10,000 for validation, and 10,000 for testing.
@@
For all models, we use a batch size of 256 with a grid sampling size of 128, except for the FNO, which uses the full grid size of 200. The models are trained using the AdamW optimizer [71] with an exponential decay learning rate scheduler. The models are trained for 200,000 iterations, and the best model state is selected based on the validation set performance to avoid overfitting. The final performance is evaluated on the test set of 10,000 samples.
@@
Due to the discontinuous nature of the problem, the relative L2 norm may not properly reflect the performance of the models. Instead, a more suitable metric for this task is the total variation [72], defined as
@@
C.4 Shallow-Water equations
@@
The dataset is generated by PDEArena [57] using SpeedyWeather.jl on a regular grid with spatial resolution of $192\times 96\left(\Delta x=1.875^{\circ},\Delta y=3.75^{\circ}\right)$, and temporal resolution of $\Delta t=48\mathrm{~{}h}$. The resulting dataset contains 6,600 trajectories with a temporal spatial resolution $11\times 96\times 192$.
@@
We follow the problem setup reported in PDEArena [57]. Our goal is to learn the operator that maps the vorticity and pressure fields from the previous 2 two time-steps to the next time-step. All models are trained with 5,600 trajectories and evaluated on the remain 1,000 trajectories. We report the relative $L^{2}$ errors over 5-steps rollout predictions.
@@
For training CViT models, please refer to section C.1. For training other baselines in Table 5, we follow the recommended training procedure and hyper-parameters presented in PDEArena [57] Specifically, we use the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay of $10^{-5}$. Training is conducted for 50 epochs, minimizing the summed mean squared error. We use cosine annealing as the learning rate scheduler [73] with a linear warm-up phase of 5 epochs. An effective batch size of 32 is used for training.
@@
C.5 Navier-Stokes equation
@@
The 2D Navier-Stokes data is generated by PDEArena [57], which is obtained on a grid with a spatial resolution of $128\times 128(\Delta x=0.25,\Delta y=0.25)$ and a temporal resolution of $\Delta t=1.5\mathrm{~{}s}$. The simulation uses a viscosity parameter of $\nu=0.01$ and a buoyancy factor of $(0,0.5)^{T}$. The equation is solved on a closed domain with Dirichlet boundary conditions $(v=0)$ for the velocity, and Neumann boundaries $\left(\frac{\partial s}{\partial x}=0\right)$ for the scalar field. The simulation runs for $21\mathrm{~{}s}$, sampling every $1.5\mathrm{~{}s}$ Each trajectory contains scalar and vector fields at 14 different time frames, resulting in a dataset with 7,900 trajectories.
@@
We follow the problem setup reported in Hao et al. [60]. We aim to learn the solution operator that maps the passive scalar and velocity fields from the previous 10 time-steps to the next time-step. The models are trained with 6,500 trajectories and tested on the remaining 1,300 trajectories. We report the resulting relative $L^{2}$ errors over 4-steps rollout predictions.
@@
We directly report the results as presented in DPOT [60]. For comprehensive details on the training procedures, hyper-parameters, please refer to their original paper.
Figure 11: Naiver-Stokes benchmark. Representative CViT rollout prediction of the passive scalar field, and point-wise error against the ground truth.
Figure 12: Naiver-Stokes benchmark. Representative CViT rollout prediction of the velocity field in $x$ direction, and point-wise error against the ground truth.
Figure 13: Naiver-Stokes benchmark. Representative CViT rollout prediction of the velocity field in $y$ direction, and point-wise error against the ground truth.
C.6 Computational Cost

### v2

@@ F.1 Training and evaluation
We use a unified training recipe for all CViT experiments. We employ AdamW optimizer (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with a weight decay $10^{-5}$. Our learning rate schedule includes an initial linear warm-up phase of $5,000$ steps, starting from zero and gradually increasing to $10^{-3}$, followed by an exponential decay at a rate of $0.9$ for every $5,000$ steps. The loss function is a one-step mean squared error (MSE) between the model predictions and the corresponding targets at the next time-step, evaluated at randomly sampled query coordinates:
@@
where $s^{(k)}_{i}(\mathbf{y_{j}})$ denotes the $k$-th variable of the $i$-th sample in the training dataset, evaluated at a query coordinate $\mathbf{y}_{j}$, and $\hat{s}$ denotes the corresponding model prediction. It is worth noting that our objective function is a one-step loss, which differs significantly from the rollout loss used in training DPOTs All models are trained for $2\times 10^{5}$ iterations with a batch size $B=64$. Within each batch, we randomly sample $Q=1,024$ query coordinates from the grid and corresponding output labels.
@@
After training, we obtain the predicted trajectory by performing an auto-regressive rollout on the test dataset. We evaluate model accuracy using the relative $L^{2}$ norm, as commonly used in Li et al.(Li et al., 2021):
@@ F.2 Model Details
Following the comprehensive experiments and ablation studies in PDEArena (Gupta & Brandstetter, 2022), we adopt the optimal Dilated ResNet configuration reported in their work. This model has 128 channels and employs group normalization. It consists of four residual blocks, each comprising seven dilated CNN layers with dilation rates of [1, 2, 4, 8, 4, 2, 1].
@@
Following the comprehensive experiments and ablation studies in PDEArena (Gupta & Brandstetter, 2022), we adopt the optimal UNO of 128 channels.
@@
For the shallow water equation, we employ the best FNO model from comprehensive experiments and ablation studies in PDEArena (Gupta & Brandstetter, 2022). Specifically, we use the FNO consisting of 4 FNO layers, each with 128 channels and 32 modes. Additionally, we use two embedding layers and two output layers with kernel sizes of $1\times 1$, as suggested in Li et al.(Li et al., 2021). We use GeLU activation functions and no normalization scheme.
For the Navier-Stokes equation, we follow the problem setup in DPOT (Hao et al., 2024), which differs from the setup in PDEArena (Gupta & Brandstetter, 2022). Therefore, we directly report the results from DPOT (Hao et al., 2024).
FFNO / GK-T / OFormer / GNOT / DPOT / MPP.
We directly report the results from DPOT (Hao et al., 2024). Detailed implementations can be found in their paper.
F.3 Linear Advection Equation
@@
We make use of the datasets released by Hoop et al.(de Hoop et al., 2022). This dataset consists of 40,000 samples of discretized initial conditions on a grid of $N=200$ points.
@@
We follow the problem setup by Hoop et al.(de Hoop et al., 2022). Our objective is to predict the solution profile at time $t=0.5$. We compare the CViT model’s performance against several strong baseline neural operators: DeepONet (Lu et al., 2021), NoMaD (Seidman et al., 2022) and Fourier Neural Operators (Li et al., 2021). For the training and evaluation, we considered a split of 20,000 samples used for training, 10,000 for validation, and 10,000 for testing.
@@
For all models, we use a batch size of 256 with a grid sampling size of 128, except for the FNO, which uses the full grid size of 200. The models are trained using the AdamW optimizer (Loshchilov & Hutter, 2017) with an exponential decay learning rate scheduler. The models are trained for 200,000 iterations, and the best model state is selected based on the validation set performance to avoid overfitting. The final performance is evaluated on the test set of 10,000 samples.
@@
Due to the discontinuous nature of the problem, the relative L2 norm may not properly reflect the performance of the models. Instead, a more suitable metric for this task is the total variation (Stein & Shakarchi, 2009), defined as
@@
F.4 Shallow-Water equations
@@
The dataset is generated by PDEArena (Gupta & Brandstetter, 2022) using SpeedyWeather.jl on a regular grid with spatial resolution of $192\times 96\left(h=1.875^{\circ},\Delta y=3.75^{\circ}\right)$, and temporal resolution of $\Delta t=48\mathrm{~{}h}$. The resulting dataset contains 6,600 trajectories with a temporal spatial resolution $11\times 96\times 192$.
@@
We follow the problem setup reported in PDEArena (Gupta & Brandstetter, 2022). Our goal is to learn the operator that maps the vorticity and pressure fields from the previous 2 two time-steps to the next time-step. All models are trained with 5,600 trajectories and evaluated on the remain 1,000 trajectories. We report the relative $L^{2}$ errors over 5-steps rollout predictions.
@@
For training CViT models, please refer to section F.1. For training other baselines in Table LABEL:tab:_swe, we follow the recommended training procedure and hyper-parameters presented in PDEArena (Gupta & Brandstetter, 2022) Specifically, we use the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay of $10^{-5}$. Training is conducted for 50 epochs, minimizing the summed mean squared error. We use cosine annealing as the learning rate scheduler (Loshchilov & Hutter, 2016) with a linear warm-up phase of 5 epochs. An effective batch size of 32 is used for training.
Impact of Grid Resolution.
In our experiments, we matched the resolution of the embedding grid to the dataset for convenience. However, our model’s performance is not significantly affected as long as the embedding grid roughly matches the dataset resolution.
To validate this, we conducted experiments with three different grid resolutions: 95×191,97×95$\times$ 191,97$\times$95 × 191 , 97 × 193 , and $128\times 200$. We performed these experiments using the shallow water equation example under the same hyper-parameter settings. The resulting relative $L^{2}$ errors are summarized below:
Table 6: Swallow Water equation: Relative $L^{2}$ error for different grid resolutions.
Grid Resolution $96\times 192$ $95\times 191$ $97\times 193$ $128\times 200$
Relative $L^{2}$ Error $1.563\times 10^{-2}$ $1.572\times 10^{-2}$ $1.568\times 10^{-2}$ $1.562\times 10^{-2}$
The results demonstrate negligible variation in error across different resolutions. This empirical evidence supports our claim that CViT’s performance is robust to small variations in grid resolution, suggesting that precise matching between the embedding grid and dataset resolution is not critical for model efficacy.
@@
F.5 Navier-Stokes equation
@@
The 2D Navier-Stokes data is generated by PDEArena (Gupta & Brandstetter, 2022), which is obtained on a grid with a spatial resolution of $128\times 128(h=0.25,\Delta y=0.25)$ and a temporal resolution of $\Delta t=1.5\mathrm{~{}s}$. The simulation uses a viscosity parameter of $\nu=0.01$ and a buoyancy factor of $(0,0.5)^{T}$. The equation is solved on a closed domain with Dirichlet boundary conditions $(v=0)$ for the velocity, and Neumann boundaries $\left(\frac{\partial s}{\partial x}=0\right)$ for the scalar field. The simulation runs for $21\mathrm{~{}s}$, sampling every $1.5\mathrm{~{}s}$ Each trajectory contains scalar and vector fields at 14 different time frames, resulting in a dataset with 7,900 trajectories.
@@
We follow the problem setup reported in Hao et al. (Hao et al., 2024). We aim to learn the solution operator that maps the passive scalar and velocity fields from the previous 10 time-steps to the next time-step. The models are trained with 6,500 trajectories and tested on the remaining 1,300 trajectories. We report the resulting relative $L^{2}$ errors over 4-steps rollout predictions.
@@
We directly report the results as presented in DPOT (Hao et al., 2024). For comprehensive details on the training procedures, hyper-parameters, please refer to their original paper.
Figure 11: Navier-Stokes benchmark. Representative CViT rollout prediction of the velocity field in $x$ direction, and point-wise error against the ground truth.
Figure 12: Navier-Stokes benchmark. Representative CViT rollout prediction of the velocity field in $y$ direction, and point-wise error against the ground truth.
F.6 Diffusion-reaction equation
The 2D diffusion-reaction equation is given by
$\partial_{t}u=D_{u}\partial_{xx}u+D_{u}\partial_{yy}u+R_{u},$
$\partial_{t}v=D_{v}\partial_{xx}v+D_{v}\partial_{yy}v+R_{v},$
where $u=u(t,x,y)$ is the the activator and $v=v(t,x,y)$ denotes the inhibitor. Here $D_{u}$ and $D_{v}$ are the diffusion coefficient for the activator and inhibitor, respectively, $R_{u}=$ $R_{u}(u,v)$ and $R_{v}=R_{v}(u,v)$ are the activator and inhibitor reaction function, respectively. The domain of the simulation includes $x\in(-1,1),y\in(-1,1),t\in(0,5]$. The reaction functions for the activator and inhibitor are defined by:
$R_{u}(u,v)=u-u^{3}-k-v,$
$R_{v}(u,v)=u-v,$
where $k=5\times 10^{-3}$, and the diffusion coefficients for the activator and inhibitor are $D_{u}=1\times 10^{-3}$ and $D_{v}=5\times 10^{-3}$, respectively. The initial condition is generated as standard normal random noise $u(0,x,y)\sim\mathcal{N}(0,1.0)$ for $x\in(-1,1)$ and $y\in(-1,1)$.
Data Generation.
We use the dataset generated by PDEBench (Takamoto et al., 2022), where discretized into $N_{x}=512,N_{y}=512$ and $N_{t}=501$, as well as the downsampled version for the models training with $N_{x}=128,N_{y}=128$, and $N_{t}=101$, the spatial discretization is performed using the finite volume method, and the time integration is performed using the built-in fourth order Runge-Kutta method in the scipy package.
Problem Setup.
We follow the problem setup reported in Hao et al. (Hao et al., 2024). We aim to learn the solution operator that maps the activator and inhibitor fields from the previous time-steps to the next time-step. The models are trained with 900 trajectories and tested on the remaining 100 trajectories. We report the resulting relative $L^{2}$ errors over rollout predictions.
Training.
For training CViT models, please refer to section F.1. It is worth noting that for this specific task, we employ an initial learning rate of $5\times 10^{-4}$, different from the $10^{-3}$ used in other experiments.
Evaluation.
We directly report the results as presented in DPOT (Hao et al., 2024). For comprehensive details on the training procedures, hyper-parameters, please refer to their original paper.
Figure 13: Diffusion-reaction Benchmark. Representative CViT rollout prediction of the density field $v$, and point-wise error against the ground truth.
F.7 Computational Cost